OpenStudy (anonymous):

find all complex numbers z such that z^2 = 7 - 24i give your answer in the form a+bi

OpenStudy (anonymous):



OpenStudy (anonymous):

(a+bi)(a+bi)=7-24i a^2+2abi-b^2=7-24i now identify real and imaginary parts: a^2-b^2=7 2ab=-24 and solve this system for a and b

OpenStudy (anonymous):

apparently im only supposed to get values of -4 and 4 for a...i get something different

OpenStudy (anonymous):

because first i isolate for b = -12/a then i get 7 = (a-(12/a))(a+(12/a))

OpenStudy (anonymous):

a=-12/b substitute in 1º equation: 144/b^2-b^2=7 144/b^2=7+b^2 144=7b^2+b^4 new variable: b^2=x 144=7x+x^2 or x^2+7x-144=0 solve this

OpenStudy (anonymous):

x = -16, 9

OpenStudy (anonymous):

b^2 = -16 b^2 = 9 cant take b^2 = -16

OpenStudy (anonymous):

so b = +/- 3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so: b^2=9 b=+-3

OpenStudy (anonymous):

ok, i got it now...plug +/- 3 in equation

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

in a=-12/b

OpenStudy (anonymous):

yw